OpenStudy (anonymous):

tan2theta - cot2theta = 0 for 0 < theta < 180degrees

OpenStudy (raden):

tan2x - cot2x = 0 tan2x - 1/tan2x = 0 multiply by tan2x to both sides, gives tan^2 2x - 1 = 0 (tan2x+1)(tan2x-1) = 0 for zeroes, take : tan2x+1=0 or tan2x-1=0 for tan2x+1=0, case I : tan2x+1=0 tan2x=-1 tan2x=tan3pi/4 tan2x=tan(k*pi+3pi/4) 2x = k*pi + 3pi/4 x = k*pi/2 + 3pi/8 case II : tan2x=-1 tan2x=tan7pi/4 but, u neednt solve it (because just for interval 0 < x < 180degrees) for tan2x-1=0 case I : tan2x-1=0 tan2x = 1 tan2x = tan pi/4 tan2x = tan (k*pi + pi/4) 2x = k*pi + pi/4 x = k*pi/2 + pi/8 case II : tan2x = 1 tan2x = tan 5pi/4 but, needn't solve it (because just for interval 0 < x < 180degrees) sorry, i was change theta be "x" :)